OpenStudy (anonymous):

the density of tin is 7.28 g/cm cubed, what is the thickness of a piece of tin foil that measures 5.70 cm by 4.25 cm and has a mass of 0.655 g

OpenStudy (aaronq):

Find the volume, \(\sf density=\dfrac{mass}{volume}\) From the volume you can get the missing dimension. \(\sf volume = l*h*w\)

OpenStudy (anonymous):

(0.655 g) / (7.28 g/cm^3) / (5.70 cm x 4.25 cm) = 0.00371 cm